tyricwoods223:

A bagel shop sells coffee in a container shaped like a rectangular prism. A graphic designer who works for the bagel shop drew the net below to create a design for the container.

Mario0120:

What's the question?

tyricwoods223:

What is the surface area of the container?

Markthegoat1117:

Is there I pic?

Mario0120:

The surface area (A) of a rectangular prism is given by the formula: A = 2lw + 2lh + 2whA=2lw+2lh+2wh where: ll is the length, ww is the width, and hh is the height. If you can provide the values for these dimensions, I can help you compute the surface area of the container.

surjithayer:

surface area of a prism=2(l+w)h+2lw =2lh+2wh+2lw

surjithayer:

correctio n rectangular prism

toga:

To find the surface area of the container, we need to find the area of each face and then add them up. Looking at the net provided, we can see that the container has six faces: a bottom, a top, a front, a back, a left side, and a right side. The bottom and top are both rectangles with dimensions of 6 inches by 4 inches. So the area of each is: 6 inches x 4 inches = 24 square inches The front and back are both rectangles with dimensions of 6 inches by 2 inches. So the area of each is: 6 inches x 2 inches = 12 square inches The left and right sides are both rectangles with dimensions of 4 inches by 2 inches. So the area of each is: 4 inches x 2 inches = 8 square inches To find the total surface area, we need to add up all six faces: Total surface area = 2(24) + 2(12) + 2(8) Total surface area = 48 + 24 + 16 Total surface area = 88 square inches Therefore, the surface area of the container is 88 square inches.

shelovejayy:

45y244 t4t42